# RIS-Empowered OTFS Modulation With Faster-than-Nyquist Signaling in High-Mobility Wireless Communications

### v1

@@ III-A AFER
$\Delta_{ij}^{\dagger}\mathbf{H}_{\mathrm{eff}}^{\dagger}\mathbf{\Sigma}_{z}^{-1}\mathbf{H}^{\mathrm{eff}}\Delta_{ij}=\operatorname{vec}(\mathbf{H}_{\mathrm{eff}})^{\dagger}\bigl(\Delta_{ij}\Delta_{ij}^{\dagger}\otimes\mathbf{\Sigma}_{z}^{-1}\bigr)\operatorname{vec}(\mathbf{H}^{\mathrm{eff}}).$ (29)
Under the assumption that $\mathbf{H}^{\mathrm{eff}}$ is composed of i.i.d. Rayleigh fading coefficients, each $\tilde{h}_{n}$ follows an independent complex Gaussian distribution with zero mean and variance $\sigma_{h}^{2}$, i.e., $\tilde{h}_{n}\sim\mathcal{CN}(0,\sigma_{h}^{2})$. Then, we can use the well-known moment-generating function (MGF) of complex Gaussian quadratic forms. For any Hermitian positive semidefinite matrix $\mathbf{A}$, we have
$\mathbb{E}\!\left[e^{-\operatorname{vec}(\mathbf{H}^{\mathrm{eff}})^{\dagger}\mathbf{A}\operatorname{vec}(\mathbf{H}^{\mathrm{eff}})}\right]=\det\!\big(\mathbf{I}+\sigma_{h}^{2}\mathbf{A}\big)^{-1}.$ (30)
@@
The AFER upper bound implicitly incorporates the transmit energy $E_{f}$ through $\sigma_{h}^{2}\lambda_{n}^{(ij)}\nu_{m}$, reflecting the SNR via the channel variance $\sigma_{h}^{2}$ and noise power $\sigma^{2}$, as embodied through the symbol difference $\Delta_{ij}g\left(n-m\right)$ and noise whitening $\mathbf{\Sigma}_{z}^{-1}$. A higher $E_{f}$ enhances signal strength, reducing the error probability, and this effect is indirectly reflected through the channel gain $\sigma_{h}^{2}$ and symbol energy distribution. Besides, due to $E_{f}=MNE_{x}$, when $M$ and $N$ increase, $E_{f}$ grows linearly, and $\text{SNR}=\frac{E_{f}}{\sigma^{2}}\propto MN$ increases. Note that $\mathbf{\Sigma}_{z}$ and $\mathbf{G}$ are positive-definite matrices. In the formula, the number of terms in $\prod_{m=1}^{MN}$ rises from $MN$ to $M^{\prime}N^{\prime}$ with more $\nu_{m}$ terms involved, leading to lower AFER. More insighful finding and insights for the AFER performance are given in Sec. V.
@@ III-B Spectral Efficiency
$=\log_{2}\det\left(\mathbf{I}_{MN}+\frac{1}{\sigma^{2}}\mathbf{G}_{\mathrm{eq}}^{-1}\mathbf{H}^{\mathrm{eff}}\mathbf{\Sigma}_{X}\mathbf{H}_{\mathrm{eff}}^{\dagger}\right).$
To simplify, perform eigenvalue decomposition (EVD) on $\mathbf{G}_{\mathrm{eq}}=\mathbf{V\Lambda V}^{H}$, where $\mathbf{V}=\mathrm{diag}\left\{\lambda^{1}_{0},\lambda^{1}_{1},...,\lambda^{1}_{M}N-1\right\}$ with $i$-th $\lambda^{1}_{i}$ non-zero eigenvalues of $\mathbf{G}_{\mathrm{eq}}$. Then, given that $\mathbf{B}=\mathbf{\Lambda}^{-1/2}\mathbf{V}^{H}\mathbf{H}^{\mathrm{eff}}$, Eq. (37) can be further rewritten as $I(\mathbf{X}_{\mathrm{DD}};\mathbf{y})=\log_{2}\det\left(\mathbf{I}_{MN}+\frac{E_{x}}{\sigma^{2}}\mathbf{B}^{H}\mathbf{B}\right)$. Let $\mathbf{B}^{H}\mathbf{B}=\mathbf{U\Xi U}^{H}$ with $\mathbf{\Xi}=\mathrm{diag}\left\{\mathrm{\xi}_{0},\mathrm{\xi}_{1},...,\mathrm{\xi}_{MN-1}\right\}$ and $\mathrm{\xi}_{i}$ being $i$-th non-zeros eigenvalue of $\mathbf{B}^{H}\mathbf{B}$. Eq. (37) can be simplified as
@@ III-C PAPR and CCDF
To obtain the CCDF curve, a large number of frames are generated using Monte Carlo simulations. For each frame, the PAPR is calculated, and the empirical distribution is estimated by counting the proportion of frames whose PAPR exceeds a given threshold $\gamma_{0}$. More results and insights are given in section Sec. V.
@@ III-D IBO
The IBO is a key parameter for assessing the linearity margin of power amplifiers (PA) and it directly relates to the peak-to-average power ratio (PAPR) of the transmit waveform. In this subsection we distinguish two useful IBO definitions used in our analysis: (i) the minimum required IBO that guarantees no clipping of waveform peaks, and (ii) the available IBO that results from a given PA saturation level and the transmitter average power. Based on Eq. (40), the required IBO for linear operation is mainly determined by the PAPR of the transmitted signal plus a design margin to avoid nonlinear distortion, expressed as
@@
IV RIS Phase Design
In our works, the goal of RIS phase optimization is to adjust the phase shifts of each RIS reflecting element to maximize the effective channel gain at the receiver. A key characteristic of an RIS system is its capability to dynamically reconfigure reflection coefficients in accordance with channel conditions. In OTFS modulation, the channel is represented in the DD domain, where variations occur at a slower rate compared to the time-frequency domain representation. Consequently, the phase adjustment frequency in the RIS can be reduced. Here, we consider the discrete phase adjustment, of which the optimization procedure is described below.
@@ IV RIS Phase Design
with $\theta_{1}=2\pi\nu_{q,p_{1,2}}^{\mathrm{eff}}(t-\tau_{q,p_{1,2}}^{\mathrm{eff}})$ and $\theta_{2}=2\pi\frac{k_{q,p_{1,2}}(a-\varepsilon_{q,p_{1,2}})}{NM}$. Detail steps for optimization can be found in Algorithm I.
@@
According to (39), the spectral efficiency is expected to improve by employing the RIS phase optimization strategy. This is because optimizing the RIS phase shifts enhances the effective channel gain by aligning the reflected signals constructively at the receiver. As a result, the matrix product $\mathbf{H}^{\mathrm{eff}}\mathbf{H}_{\mathrm{eff}}^{H}$ contains stronger eigenvalues, leading to a larger determinant inside the logarithm. Although random phase adjustment may also increase compared to the case without RIS, which is because random phase configurations can still introduce diversity and enable some constructive interference at the receiver by chance, the effective channel gain $\mathbf{H}_{\mathrm{eff}}$ is generally weaker than in the case of optimal phase alignment since the phase shifts are not aligned to maximize the received signal power Consequently, the performance gain from random phase modulation is typically limited and less consistent. More corresponding results and insights are given in Sec. V.
@@
Figure 5: IBO of the RIS-OTFS-FTN and OTFS-FTN schemes with $\alpha=0.8$, $M=32$ and $N=32$.
@@
Figure 7: BER of the MMSE detector-assisted RIS-OTFS-FTN scheme with various comparing cases: (a) $Q=4$ and $\alpha=0.7,0.8,0.9$; (b) $\alpha=0.8$ and $Q=8,16,32$; (c) $\alpha=0.8$, $Q=8,16,32$ in optimal phase adjustment and the random phaseone (Ran.).
@@ V Simulation Results
In this section, theoretical and simulation results of the RIS-OTFS-FTN scheme, e.g., PAPR, CCDF, IBO, FER, BER, and spectral efficiency are provided. Here, to approach the channel environment in suburban or urban high-speed scenarios, a standard EVA wireless channel model is employed in this section  [^2] EVA channel model is a representative wireless fading model widely employed for simulating high-mobility environments, such as those encountered in LTE, 5G, and performance evaluations of advanced schemes like the RIS-OTFS-FTN system. As an extended version of the Vehicular A (VA) model recommended by the 3GPP and ITU standards [35], the EVA model features a longer delay spread and more complex multi-path characteristics. These enhancements allow it to more accurately capture the impact of multi-path interference in high-speed scenarios.. The detail values of simulation parameter in EVA model are given in Table I. For a clear presentation and comparison in this section, several other parameters are also fixed, including size of QAM $M_{QAM}=2$, carrier frequency $f_{c}=4$ GHz, subcarrier spacing $\bigtriangleup f=15$ KHz, unit transmitted energy of each symbol $E_{x}=1$, maximum speed for user $v_{max}=120$ km/h, and SNR as $\frac{{}_{E_{x}}}{{}^{\sigma^{2}}}$. Two parameters of RRC filter are firstly given, e.g., RRC roll-off factor $\beta_{RRC}=0.25$ and symbol period span $\mathrm{s}_{RRC}=8$. Moreover, Monte Carlo simulations are conducted with each experiment performed over $3\times 10^{3}$ independent channel realizations, while considering the large complexity in calculation and running simulation. In consideration of the practical implementation of an RIS, discrete phase shifts with 3 bits allocated to each are adopted for simulation purposes, and the set of feasible values is expressed as $\theta_{q}^{v}\in\left\{0,\frac{2\pi}{8},\frac{4\pi}{8},\frac{6\pi}{8},\pi,\frac{10\pi}{8},\frac{12\pi}{8},\frac{14\pi}{8}\right\}$. The results are presented under the assumption that there exists no direct vehicle-user channel, with the RI positioned equidistantly from both the vehicle and the user.
@@
Fig. 5 illustrates the input back-off (IBO) performance of the proposed RIS-OTFS-FTN scheme compared to that of the OTFS-FTN scheme without RIS assistance. Notably, considering normal configuration in the RIS and better presentation, a fixed value of $G_{RIS}=6$ dB is given here. Here, the available IBO is computed according to (43), where the PA saturation power is fixed and the average transmit power is determined based on $E_{f}$. As shown in the figure, the RIS-assisted scheme consistently achieves a higher available IBO across the entire SNR range, with the improvement becoming more pronounced at higher SNR values. This behavior is attributed to the additional channel gain introduced by the RIS, which enables the transmitter to operate at a lower average power for the same error performance, thereby increasing the linear margin 333linear margin refers to the margin of input signal dynamic range that a system or device can withstand while maintaining the linear characteristics of the output signal. To put it simply, it is the allowable fluctuation range of input power without causing significant nonlinear distortions (such as harmonics and inter-modulation interference). A larger margin indicates a stronger ability of the system to resist distortion when dealing with signal changes. between the average transmit power and the PA saturation limit. The results clearly demonstrate the energy-efficiency benefits of incorporating RIS in high-mobility OTFS-FTN systems. In particular, the increased available IBO relaxes the linearity requirements of the PA, reduces the risk of nonlinear distortion, and allows for more efficient operation without compromising signal integrity. From a system design perspective, this implies that the RIS-OTFS-FTN scheme can maintain robust performance even when operating closer to the PA’s nonlinear region, thus reducing hardware constraints and improving overall power efficiency. Moreover, the observed IBO advantage complements the PAPR analysis in Fig. 4, as the RIS gain not only offsets the minimal PAPR increase due to FTN but also extends the operational flexibility of the transmitter in practical deployments.
@@
Fig. 7 illustrates the BER performance of the MMSE detector-assisted RIS-OTFS-FTN scheme across same values of $M=16$ and $N=16$ and various configurations in $Q$ and $\alpha$, as given in title. Note that the Opt. and Ran. in legends represent the optimal RIS phase design in this paper and random phase selection, respectively. In (a), with a fixed number of RIS elements $L=4$ and different $\alpha=0.7,0.8,0.9$, the BER decreases as $\alpha$ increases, approaching the Nyquist signaling regime ($\alpha=1$). This is also because higher $\alpha$ values reduce the severity of intentional ISI introduced by FTN, leading to improved detection accuracy. However, the differences in BER among these $\alpha$ values are relatively approached, particularly at low-to-moderate SNR, indicating that the spectral efficiency gains from lower $\alpha$ come at a limited cost in error performance under the MMSE detector. (b) examines the impact of increasing reflection elements in the RIS ($L=8,16,32$) with fixed $\alpha=0.8$. As $Q$ increasing, the BER performance improves significantly due to enhanced channel gain through the RIS. Notably, an error floor trend emerges at higher SNR, becoming more pronounced and appearing earlier at lower SNR with larger $Q$. This behavior is attributed to the increased ISI in the cascaded vehicle-RIS-user channels, where more reflection elements amplify multi-path components, exacerbating interference in the DD domain and limiting the MMSE detector’s ability to fully mitigate it. For (c) ,the Ran. design yields substantially higher BER across all SNR by comparing to the Opt. design, with differences exceeding an order of magnitude in some regions. This underscores the critical role of Algorithm 1, which aligns reflected signals constructively to maximize channel gain.
@@
Fig. 8 depicts the BER performance of the proposed RIS-OTFS-FTN scheme with MMSE detector in comparison to benchmark variants, including schemes without RIS or FTN, as well as different $\alpha$ $M$ and $N$., as shown in figure. Notably, the RIS-OTFS scheme ($\alpha=1$) achieves superior performance relative to FTN-enabled variants at equivalent grid sizes, attributable to its use of ideal rectangular pulses rather than the RRC filter employed in FTN processing. Moreover, the RIS-OTFS-FTN with $M=16$, $N=16$, $Q=8$, and $\alpha=0.8$ extremely outperforms the OTFS-FTN without RIS, demonstrating that RIS can better enhance the effective channel strength and mitigates fading in high-mobility scenarios. Similarly, comparing RIS-OTFS-FTN to RIS-OTFS, the FTN variant introduces additional ISI due to time-domain compression, leading to a performance degradation, particularly evident at higher SNR where an error floor emerges, which is because increased grid dimensions amplify the impact of residual ISI and channel dispersion in the DD domain, limiting the MMSE detector’s equalization capability. Besides, we can find that the BER performance improves as $M$ and $N$ increases, which happens because larger OTFS grids offer finer resolution in the DD domain, enabling better exploitation of channel diversity and more effective equalization of multi-path and Doppler effects in detection.

### v2

@@ III-A AFER
$\Delta_{ij}^{\dagger}\mathbf{H}_{\mathrm{eff}}^{\dagger}\mathbf{\Sigma}_{z}^{-1}\mathbf{H}_{\mathrm{eff}}\Delta_{ij}=\operatorname{vec}(\mathbf{H}_{\mathrm{eff}})^{\dagger}\bigl(\Delta_{ij}\Delta_{ij}^{\dagger}\otimes\mathbf{\Sigma}_{z}^{-1}\bigr)\operatorname{vec}(\mathbf{H}_{\mathrm{eff}}).$ (29)
WIth assumption that $\mathbf{H}^{\mathrm{eff}}$ is composed of i.i.d. Rayleigh fading coefficients, each $\tilde{h}_{n}$ follows an independent Gaussian distribution with zero mean and variance $\sigma_{h}^{2}$, i.e., $\tilde{h}_{n}\sim\mathcal{CN}(0,\sigma_{h}^{2})$. We can use the well-known moment-generating function (MGF) of complex Gaussian quadratic forms. For any Hermitian positive semidefinite matrix $\mathbf{A}$, we have
$\mathbb{E}\!\left[e^{-\operatorname{vec}(\mathbf{H}_{\mathrm{eff}})^{\dagger}\mathbf{A}\operatorname{vec}(\mathbf{H}_{\mathrm{eff}})}\right]=\det\!\big(\mathbf{I}+\sigma_{h}^{2}\mathbf{A}\big)^{-1}.$ (30)
@@
A higher $E_{f}$ enhances signal strength, reducing the error probability, and this effect is indirectly reflected through the channel gain and symbol energy distribution. Besides, due to $E_{f}=MNE_{x}$, when $M$ and $N$ increase, $E_{f}$ grows linearly, and $\text{SNR}=\frac{E_{f}}{\sigma^{2}}\propto MN$ increases. Note that $\mathbf{\Sigma}_{z}$ and $\mathbf{G}$ are positive-definite matrices. In the formula, the number of terms in $\prod_{m=1}^{MN}$ rises from $MN$ to $M^{\prime}N^{\prime}$ with more $\nu_{m}$ terms involved, leading to lower AFER.
Remark 2
From the eigenvalue decomposition of the PEP upper bound in (28), let $r_{ij}\triangleq\mathrm{rank}(\boldsymbol{\Delta}_{ij}\boldsymbol{\Delta}_{ij}^{\dagger})=\mathrm{rank}(\boldsymbol{\Delta}_{ij})$ denote the number of non-zero eigenvalues of the error difference matrix $\boldsymbol{\Delta}_{ij}\boldsymbol{\Delta}_{ij}^{\dagger}$. Given that the FTN-induced noise covariance $\mathbf{\Sigma}_{z}=\sigma^{2}\mathbf{G}(\alpha)$ is positive definite, its inverse eigenvalues scale as $\nu_{m}(\alpha)=\mu_{m}(\alpha)/\sigma^{2}$, where $\{\mu_{m}(\alpha)\}$ are the eigenvalues of $\mathbf{G}^{-1}(\alpha)$ which are independent of the noise power $\sigma^{2}$. In the high-SNR regime, each term in the product expansion behaves as $(1+c\lambda_{n}^{(ij)}\nu_{m}(\alpha))^{-1}\doteq\mathrm{SNR}^{-1}$. This yields the asymptotic decay rate of the error probability as
$P\!\left(\mathbf{x}^{\mathrm{DD}}_{i}\!\to\!\mathbf{x}^{\mathrm{DD}}_{j}\right)\doteq\mathrm{SNR}^{-MN\cdot r_{ij}}.$ (34)
Thus, asymptotic diversity order is given by $d=MN\cdot\min_{i\neq j}r_{ij}$. Crucially, the FTN acceleration factor $\alpha$ modifies the eigen-spectrum $\{\mu_{m}(\alpha)\}$ of the correlation matrix $\mathbf{G}(\alpha)$, thereby affecting coding gain. However, it does not alter the diversity exponent (slope) as long as $\mathbf{G}(\alpha)$ remains full rank.
@@ III-B Spectral Efficiency
$=\log_{2}\det\left(\mathbf{I}_{MN}+\frac{1}{\sigma^{2}}\mathbf{G}^{-1}\mathbf{H}^{\mathrm{eff}}\mathbf{\Sigma}_{X}\mathbf{H}_{\mathrm{eff}}^{\dagger}\right).$
To simplify, perform eigenvalue decomposition (EVD) on $\mathbf{G}=\mathbf{V\Lambda V}^{H}$, where $\mathbf{V}=\mathrm{diag}\left\{\lambda^{1}_{0},\lambda^{1}_{1},...,\lambda^{1}_{M}N-1\right\}$ with $i$-th $\lambda^{1}_{i}$ non-zero eigenvalues of $\mathbf{G}$. Then, given that $\mathbf{B}=\mathbf{\Lambda}^{-1/2}\mathbf{V}^{H}\mathbf{H}^{\mathrm{eff}}$, Eq. (38) can be further rewritten as $I(\mathbf{X}_{\mathrm{DD}};\mathbf{y})=\log_{2}\det\left(\mathbf{I}_{MN}+\frac{E_{x}}{\sigma^{2}}\mathbf{B}^{H}\mathbf{B}\right)$. Let $\mathbf{B}^{H}\mathbf{B}=\mathbf{U\Xi U}^{H}$ with $\mathbf{\Xi}=\mathrm{diag}\left\{\mathrm{\xi}_{0},\mathrm{\xi}_{1},...,\mathrm{\xi}_{MN-1}\right\}$ and $\mathrm{\xi}_{i}$ being $i$-th non-zeros eigenvalue of $\mathbf{B}^{H}\mathbf{B}$. Eq. (38) can be simplified as
@@ III-C PAPR and CCDF
To obtain the CCDF curve, a large number of frames are generated using Monte Carlo simulations. For each frame, the PAPR is calculated, and the empirical distribution is estimated by counting the proportion of frames whose PAPR exceeds a given threshold $\gamma_{0}$.
@@ III-D IBO
The IBO is a key parameter for assessing the linearity margin of PA and it directly relates to the PAPR of the transmit waveform. In this subsection we distinguish two useful IBO definitions used in our analysis: (i) the minimum required IBO that guarantees no clipping of waveform peaks, and (ii) the available IBO that results from a given PA saturation level and the transmitter average power. Based on Eq. (41), the required IBO for linear operation is mainly determined by the PAPR of the transmitted signal plus a design margin to avoid nonlinear distortion, expressed as
@@
III-E Complexity Analysis and Scalability
The introduction of FTN induces colored noise and intentional ISI, creating a coupled effective channel structure. However, since the pulse shaping is localized, the FTN processing involves banded convolution scaling linearly as $\mathcal{O}(MNW)$, where $W$ is the effective pulse span. For the RIS configuration, the proposed phase alignment, e.g., Algorithm 1 in Sec. IV, incurs a low complexity of $\mathcal{O}(PQ)$. Crucially, this overhead is amortized over the channel coherence time, as RIS phases are not updated on a per-symbol basis. Regarding detection, while the adopted LMMSE detector theoretically entails $\mathcal{O}((MN)^{3})$ complexity via direct inversion, the effective channel matrix retains a banded-block-Toeplitz structure. For practical scalability, this structure can be exploited by advanced banded solvers to reduce the complexity to approximately $\mathcal{O}(N_{\mathrm{sys}}W^{2})$, where $N_{\mathrm{sys}}\approx MN$ denotes the dimension of the underlying linear system, i.e., the number of DD symbols per OTFS frame. This ensures that the proposed scheme remains tractable for moderate grid sizes while offering significant spectral efficiency gains.
@@ III-E Complexity Analysis and Scalability
IV RIS Design
IV-A Adjusted Phase Algorithm
In our work, the goal of RIS phase optimization is to adjust the phase shifts of each RIS reflecting element to maximize the effective channel gain at the receiver. A key characteristic of an RIS system is its capability to dynamically reconfigure reflection coefficients in accordance with channel conditions. In OTFS modulation, the channel is represented in the DD domain, where variations occur at a slower rate compared to the time-frequency domain representation. Consequently, the phase adjustment frequency in the RIS can be reduced. Here, we consider the discrete phase adjustment, of which the optimization procedure is described below.
@@ IV-A Adjusted Phase Algorithm
with $\theta_{1}=2\pi\nu_{q,p_{1,2}}^{\mathrm{eff}}(t-\tau_{q,p_{1,2}}^{\mathrm{eff}})$ and $\theta_{2}=2\pi\frac{k_{q,p_{1,2}}(a-\varepsilon_{q,p_{1,2}})}{NM}$. Detail steps for optimization can be found in Algorithm 1.
@@
According to (40), the spectral efficiency is expected to improve by employing the RIS phase optimization strategy. This is because optimizing the RIS phase shifts enhances the effective channel gain by aligning the reflected signals constructively at the receiver. Thus, the matrix product $\mathbf{H}_{\mathrm{eff}}\mathbf{H}_{\mathrm{eff}}^{H}$ contains stronger eigenvalues, leading to a larger determinant inside the logarithm. Although random phase adjustment may also increase compared to the case without RIS, which is because random phase configurations can still introduce diversity and enable some constructive interference at the receiver by chance, the effective channel gain $\mathbf{H}_{\mathrm{eff}}$ is generally weaker than in the case of optimal phase alignment since the phase shifts are not aligned to maximize the received signal power Consequently, the performance gain from random phase modulation is typically limited and less consistent. More corresponding results and insights are given in Sec. V.
Remark 5
Identifying the strongest path from $P=P_{1}+P_{2}$ dominant multipath components require $\mathcal{O}(P)$ operations. Subsequently, calculating the optimal phase shifts for $Q$ reflection elements involve $\mathcal{O}(Q)$ complex additions and phase extractions. Therefore, the total computational complexity is in the order of $\mathcal{O}(P+Q)$. Compared to the signal detection complexity, e.g., $\mathcal{O}((NM)^{3})$ for LMMSE, the overhead introduced by the RIS phase design can be negligible.
IV-B Robustness to Channel Estimation Errors
In practical high-mobility scenarios with RIS applications, obtaining perfect channel state information (CSI) is challenging. To evaluate the robustness of the proposed RIS phase alignment strategy, we model the impact of imperfect CSI. For better presentation, we assume the transmitter-to-RIS link, being stationary, is accurately estimated. However, the RIS-to-User link suffers from estimation errors due to user mobility and difficulty of estimating reflecting link. The estimated channel impulse response with imperfect CSI for the RIS-to-User link is modeled as
$\hat{\boldsymbol{h}}_{q}^{2}=\boldsymbol{h}_{q}^{2}+\mathbf{e}_{q},$ (51)
where $\boldsymbol{h}_{q}^{2}$ denotes the vector of path gains $h_{q,0}^{2},\dots,h_{q,P_{2}-1}^{2}]^{T}$ corresponding to the time-domain response in (II-B), and $\mathbf{e}_{q}\sim\mathcal{CN}(\mathbf{0},\delta_{e^{2}}^{2}\mathbf{I})$ represents the channel estimation error vector.
IV-C Impact of Finite Resolution Phase Quantization
Practical RIS implementations are constrained by hardware limitations, typically supporting only a finite set of discrete phase levels. To evaluate the impact of this practical constraint, we consider a $B$-bit quantization scheme. The set of feasible discrete phase shifts is defined as $\mathcal{F}_{B}=\{0,\frac{2\pi}{2^{B}},\dots,\frac{2\pi(2^{B}-1)}{2^{B}}\}$. Consequently, the continuous optimal phase $\theta_{q}^{\mathrm{opt}}$ obtained from Algorithm 1 is quantized to the nearest discrete level $\bar{\theta}_{q}\in\mathcal{F}_{B}$ . The resulting quantized phase shift matrix, denoted as $\bar{\mathbf{\Phi}}=\mathrm{diag}(\beta_{1}e^{j\bar{\theta}_{1}},\dots,\beta_{Q}e^{j\bar{\theta}_{Q}})\otimes\mathbf{I}_{NM}$, is then applied to the effective channel construction. This quantization introduces a phase mismatch that degrades the passive beamforming gain. The trade-off between the quantization resolution $B$, e.g., 1, 2, and 3-bit, and the system error performance is numerically investigated in Section V.
@@ IV-C Impact of Finite Resolution Phase Quantization
Figure 5: IBO of the RIS-OTFS-FTN and OTFS-FTN schemes with $\alpha=0.8$ and $1$, $M=32$ and $N=32$.
@@
Figure 7: BER of the MMSE detector-assisted RIS-OTFS-FTN scheme with various comparing cases: (a) $Q=4$ and $\alpha=0.7,0.8,0.9$; (b) $\alpha=0.8$ and $Q=8,16,32$; (c) $\alpha=0.8$, $Q=8,16,32$ in optimal phase adjustment (Opt.) and the random phase one (Ran.).
@@ V Simulation Results
In this section, theoretical and simulation results of the RIS-OTFS-FTN scheme, e.g., PAPR, CCDF, IBO, FER, BER, and spectral efficiency are provided. Here, to approach the channel environment in suburban or urban high-speed scenarios, a standard EVA wireless channel model is employed in this section  [^4] EVA channel model is a representative wireless fading model widely employed for simulating high-mobility environments, such as those encountered in LTE, 5G, and performance evaluations of advanced schemes like the RIS-OTFS-FTN system. As an extended version of the Vehicular A (VA) model recommended by the 3GPP and ITU standards [36], the EVA model features a longer delay spread and more complex multi-path characteristics. These enhancements allow it to more accurately capture the impact of multi-path interference in high-speed scenarios.. The detail values of simulation parameter in EVA model are given in Table I. For a clear presentation and comparison in this section, several other parameters are also fixed, including size of QAM $M_{QAM}=2$, carrier frequency $f_{c}=4$ GHz, subcarrier spacing $\bigtriangleup f=15$ KHz, unit transmitted energy of each symbol $E_{x}=1$, maximum speed for user $v_{max}=120$ km/h, and SNR as $\frac{{}_{E_{x}}}{{}^{\sigma^{2}}}$. Two parameters of RRC filter are firstly given, e.g., RRC roll-off factor $\beta_{RRC}=0.25$ and symbol period span $\mathrm{s}_{RRC}=8$. Moreover, Monte Carlo simulations are conducted with each experiment performed over $3\times 10^{3}$ independent channel realizations, while considering the large complexity in calculation and running simulation. In consideration of the practical implementation of an RIS, discrete phase shifts with 3 bits allocated to each are adopted for simulation purposes, and the set of feasible values is expressed as $\theta_{q}^{v}\in\left\{0,\frac{2\pi}{8},\frac{4\pi}{8},\frac{6\pi}{8},\pi,\frac{10\pi}{8},\frac{12\pi}{8},\frac{14\pi}{8}\right\}$. The results are presented under the assumption that there exists no direct vehicle-user channel, with the RIS positioned equidistantly from both the vehicle and the user.
@@
Fig. 5 illustrates the input back-off (IBO) performance of the proposed RIS-OTFS-FTN scheme compared to that of the OTFS-FTN scheme without RIS assistance. Notably, considering normal configuration in the RIS and better presentation, a fixed value of $G_{RIS}=6$ dB is given here. Here, the available IBO is computed according to (44), where the PA saturation power is fixed and the average transmit power is determined based on $E_{f}$. As shown in the figure, the RIS-assisted scheme consistently achieves a higher available IBO across the entire SNR range, with the improvement becoming more pronounced at higher SNR values. This behavior is attributed to the additional channel gain introduced by the RIS, which enables the transmitter to operate at a lower average power for the same error performance, thereby increasing the linear margin 555linear margin refers to the margin of input signal dynamic range that a system or device can withstand while maintaining the linear characteristics of the output signal. To put it simply, it is the allowable fluctuation range of input power without causing significant nonlinear distortions (such as harmonics and inter-modulation interference). A larger margin indicates a stronger ability of the system to resist distortion when dealing with signal changes. between the average transmit power and the PA saturation limit. The results clearly demonstrate the energy-efficiency benefits of incorporating RIS in high-mobility OTFS-FTN systems. In particular, the increased available IBO relaxes the linearity requirements of the PA, reduces the risk of nonlinear distortion, and allows for more efficient operation without compromising signal integrity. From a system design perspective, this implies that the RIS-OTFS-FTN scheme can maintain robust performance even when operating closer to the PA’s nonlinear region, thus reducing hardware constraints and improving overall power efficiency. Moreover, the observed IBO advantage complements the PAPR analysis in Fig. 4, as the RIS gain not only offsets the minimal PAPR increase due to FTN but also extends the operational flexibility of the transmitter in practical deployments. Moreover, by comparing the results with Nyquist signaling baseline ($\alpha=1$, i.e., without FTN) alongside the proposed FTN scheme ($\alpha=0.8$), it is observed that the curve for FTN ($\alpha=0.8$) is comparable to, and in some regions even marginally lower than, that of the Nyquist baseline. That is, the proposed scheme achieves higher spectral efficiency without incurring a penalty in terms of PAPR or linearity requirements.
@@
Fig. 7 illustrates the BER performance of the MMSE detector-assisted RIS-OTFS-FTN scheme across same values of $M=16$ and $N=16$ and various configurations in $Q$ and $\alpha$, as given in title. Note that the Opt. and Ran. in legends represent the optimal RIS phase design in this paper and random phase selection, respectively. In (a), with a fixed number of RIS elements $L=4$ and different $\alpha=0.7,0.8,0.9$, the BER decreases as $\alpha$ increases, approaching the Nyquist signaling regime ($\alpha=1$). This is also because higher $\alpha$ values reduce the severity of intentional ISI introduced by FTN, leading to improved detection accuracy. However, the differences in BER among these $\alpha$ values are relatively approached, particularly at low-to-moderate SNR, indicating that the spectral efficiency gains from lower $\alpha$ come at a limited cost in error performance under the MMSE detector. (b) examines the impact of increasing reflection elements in the RIS ($L=8,16,32$) with fixed $\alpha=0.8$. As $Q$ increasing, the BER performance improves significantly due to enhanced channel gain through the RIS. Notably, an error floor trend emerges at higher SNR, becoming more pronounced and appearing earlier at lower SNR with larger $Q$. This behavior is attributed to the increased ISI in the cascaded vehicle-RIS-user channels, where more reflection elements amplify multi-path components, exacerbating interference in the DD domain and limiting the MMSE detector’s ability to fully mitigate it. For (c), the Ran. design yields substantially higher BER across all SNR by comparing to the Opt. design with differences exceeding an order of magnitude in some regions, especially with high $Q$ and in low SNR range. This underscores the critical role of Algorithm 1, which aligns reflected signals constructively to maximize channel gain.
@@
Fig. 8 depicts the BER performance of the RIS-OTFS-FTN scheme with MMSE detector in comparison to benchmark variants, including schemes without RIS or FTN, as well as different $\alpha$ $M$ and $N$., as shown in figure. Notably, the RIS-OTFS scheme ($\alpha=1$) achieves superior performance relative to FTN-enabled variants at equivalent grid sizes, attributable to its use of ideal rectangular pulses rather than the RRC filter employed in FTN processing. And the RIS-OTFS-FTN with $M=16$, $N=16$, $Q=8$, and $\alpha=0.8$ extremely outperforms the OTFS-FTN without RIS, demonstrating that RIS can better enhance the effective channel strength and mitigates fading in high-mobility scenarios. Similarly, comparing RIS-OTFS-FTN to RIS-OTFS, FTN variant introduces additional ISI due to time-domain compression, leading to a performance degradation, particularly evident at higher SNR where an error floor emerges, which is because increased grid dimensions amplify the impact of residual ISI and channel dispersion in the DD domain, limiting MMSE detector’s equalization capability. Besides, the BER performance improves as $M$ and $N$ increases, which happens because larger OTFS grids offer finer resolution in the DD domain, enabling better exploitation of channel diversity and more effective equalization of multi-path and Doppler effects in detection. For the results with message passing (MP) detector, which has 100 maximum number of MP detector iterations and damping parameter $\omega_{MP}=0.8$, we observe that despite the dense channel graph, the MP detector’s ability to explicitly model the FTN interference structure allows it to outperform LMMSE in the high-SNR regime, effectively improving the error performance in high SNR range. This confirms that while MP is computationally expensive in RIS environments, it is necessary for maximizing reliability when aggressive FTN compression is applied.